Which has more momentum?
  • A car of mass 900 kg traveling at 30 m/s
  • A railroad car of mass 20000 kg traveling at 3 m/s
  • A truck of mass 1000 kg traveling at 20 m/s 
  • A bullet of mass 0.2 kg traveling at 300 m/s
Answer #1

momentum

car = 900* 30 = 27000 kg m/s

railroad = 3* 20000 = 60000 kgm/s

truck = 1000* 20 = 20000 kgm/s

bullet = 0.2* 300 = 60 kg m/s

so railroad has highest momentum.